OpenStudy (anonymous):

The curve y = ax^3 - 24x + b has a locale minimum at (2;-17) 9.1) Calculate the values of a and b 9.2) The co-ordinate of the other turning point , E , on the curve 9.3) For what value of k wil ax^3 - 24x + b = k have three different roots. 9.4) Determine the co-ordinate of the inflection point of the curve

OpenStudy (whpalmer4):

Do you have any idea about how to proceed?

OpenStudy (anonymous):

I can do everything except 9.3.....

OpenStudy (whpalmer4):

Great! What have you found for a and b?

OpenStudy (anonymous):

My book just gives the answer k is an element of (-17;47)

OpenStudy (anonymous):

The answer to a and b appearently is a = 2 and b = 15

OpenStudy (whpalmer4):

Well, if we differentiate \(y = ax^3-24x+b\) we get \[\frac{dy}{dx} = 3ax^2-24\] and differentiating again get \(6ax\). Setting the first derivative equal to 0 and plugging in our known value of x, \[0 = 3a(2)^2-24\]\[24 = 12a\]\[a=2\]and we can find \(b\) by using the original equation and (2,-17). But you knew that... I'm a little confused by part c, I must admit, because it looks to me like there are many values of k that give 3 different roots...

OpenStudy (whpalmer4):

Oh, I didn't see that you had posted the answer. Yes, those two values make you get only two distinct roots. But how to find them...hmm...

OpenStudy (anonymous):

I dont understand it at all, I did "mistype" though it says, for what "values" not value...

OpenStudy (anonymous):

Hahaha... I should note that if you calculate the maximum and the minumum... you get the x y value of the minumum and the y value of the maximum to be respectively -17 and 47..... I still dont see the relation though....

OpenStudy (whpalmer4):

Ah, duh, of course! At those spots you have an inflection point, and that means a repeated root.

OpenStudy (whpalmer4):

I was looking at a much more difficult approach :-) If we solve \[2x^3-24x+15 == k\] the roots would be: \[\left\{\left\{x\to \frac{4\ 2^{2/3}}{\left(-15+k+\sqrt{-799-30 k+k^2}\right)^{1/3}}+\frac{\left(-15+k+\sqrt{-799-30 k+k^2}\right)^{1/3}}{2^{2/3}}\right\},\left\{x\to -\frac{2\ 2^{2/3} \left(1+i \sqrt{3}\right)}{\left(-15+k+\sqrt{-799-30 k+k^2}\right)^{1/3}}-\frac{\left(1-i \sqrt{3}\right) \left(-15+k+\sqrt{-799-30 k+k^2}\right)^{1/3}}{2\ 2^{2/3}}\right\},\left\{x\to -\frac{2\ 2^{2/3} \left(1-i \sqrt{3}\right)}{\left(-15+k+\sqrt{-799-30 k+k^2}\right)^{1/3}}-\frac{\left(1+i \sqrt{3}\right) \left(-15+k+\sqrt{-799-30 k+k^2}\right)^{1/3}}{2\ 2^{2/3}}\right\}\right\}\] (plus two more that aren't coming over) and trying to eyeball how we would make two of the three be equal was giving me quite a headache!

OpenStudy (whpalmer4):

Root 1: \[\frac{4\ 2^{2/3}}{\left(-15+k+\sqrt{-799-30 k+k^2}\right)^{1/3}}+\frac{\left(-15+k+\sqrt{-799-30 k+k^2}\right)^{1/3}}{2^{2/3}}\] Root 2: \[-\frac{2\ 2^{2/3} \left(1+i \sqrt{3}\right)}{\left(-15+k+\sqrt{-799-30 k+k^2}\right)^{1/3}}-\frac{\left(1-i \sqrt{3}\right) \left(-15+k+\sqrt{-799-30 k+k^2}\right)^{1/3}}{2\ 2^{2/3}}\] Root 3: \[-\frac{2\ 2^{2/3} \left(1-i \sqrt{3}\right)}{\left(-15+k+\sqrt{-799-30 k+k^2}\right)^{1/3}}-\frac{\left(1+i \sqrt{3}\right) \left(-15+k+\sqrt{-799-30 k+k^2}\right)^{1/3}}{2\ 2^{2/3}}\]

OpenStudy (anonymous):

wow.... Is that a formula?..

OpenStudy (whpalmer4):

If we set k = 47, they evaluate to 4, -2, and -2. If we set k = -17, they evaluate to 2, -4, and 2. Anyone can see that ;-)

OpenStudy (whpalmer4):

So, yeah, the local minima or maxima was the key. When a polynomial has an even number of roots at the same point, it is either a local minima or maxima. If it has an odd number, then it is an inflection point.

OpenStudy (whpalmer4):

Here are some graphs. No interaction with me is complete without some graph :-) You can see that one graph (k = -17) has the curve dipping down and touching the x-axis at x = 2, so that's our repeated root. The other graph (k = 47) has the curve rising up and touching the x-axis at x = -2 for our other repeated root.

OpenStudy (whpalmer4):

And for completeness, here's another graph where we don't have a duplicate root, and you see the curve crossing the x-axis 3 times: